# A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models

### v2

@@ 3.7 Generalization to Different Sizes
Next, we expand our analysis to different sizes of GPT-2 (small, large and XL). Here, the intervention results show similar suppression mechanism patterns and information flow across all models. However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix K for additional details). Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex. We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 6(a)–(b)).
3.8 Generalization to Different Models
@@
We further extend our investigation to assess whether a similar reasoning mechanism can be identified across different LMs. Specifically, we evaluate the circuit on: Pythia with 70M, 160M, 410M, and 1B parameters (Biderman et al., 2023); Qwen 2.5 with 0.5B and 1.5B parameters (Yang et al., 2024); Llama 3.2 with 1B parameters (Grattafiori et al., 2024); and additionally, GPT-2-medium after being fine-tuned on argumentative texts for critical thinking (Betz et al., 2021).
@@
Our results reveal distinct grouping patterns. In relatively small models—such as Pythia 70M and 160M—we observe erratic activation patterns that mirror the limitations seen in under-scaled models like GPT-2 small. In contrast, models with more than 0.5B parameters (namely, Pythia 410M, Pythia 1B, and Llama 3.2 1B) exhibit a compatible pattern in terms of information flow and suppressive mechanisms. Qwen exhibits a variant suppressive mechanism, with the suppression head operating at the final token position rather than $[p]$. Meanwhile, the fine-tuned GPT-2 medium model yields activation patterns that are nearly indistinguishable from its pre-trained counterpart, suggesting that the core reasoning circuit originates from pre-training rather than fine-tuning. Overall, our results indicate that while the core reasoning circuit generalizes effectively across sufficiently scaled models, subtle architectural nuances can lead to distinct but comparable internal dynamics (See Appendix L for the result details).
@@ 5 Conclusion
In this study, we have provided a comprehensive mechanistic interpretation of syllogistic reasoning within auto-regressive language models. Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms but that, at the same time, such mechanisms might be contaminated and suppressed by the same specific world knowledge acquired during pre-training.
@@ 6 Limitations
It is important to acknowledge some of the limitations of our study. First, our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios which are hard to model via causal intervention techniques. Second, Our experimentation is confined to specific syllogistic reasoning templates due to the intricate level of granularity and the computational complexity required in circuit analysis. Despite these limitations, however, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models for reasoning along with existing biases and lay a solid foundation for future research in this field.
@@ Appendix A Unconditionally Valid Syllogism Schemes
Table 3: 15 Unconditionally valid syllogism schemes. The table lists the syllogisms by their traditional names, moods, and figures, with formalised logical expressions on the first line and corresponding natural language examples on the second line. The minor premise (m) is presented before the major premise (M) to emphasise the transitive property in AAA-1 syllogism.
@@ B.2 Non-symbolic
  • Constraints: We exclude terms tokenised into multiple tokens to maintain consistency in comparison with the symbolic dataset, which is essential for coherent circuit analysis.
  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al., 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent. This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
@@ B.3 Data Statistics
We organise data statistics for generated datasets (Table 4).
@@ Appendix C Embedding Space Analysis Details
One established method for analysing hidden activation in transformer language models is the logit lens, which projects activation into embedding space (Nostalgebraist, 2020; Elhage et al., 2021; Geva et al., 2022). We focus on an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024), utilising both unembedding ($W_{U}$) and embedding ($W_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ denotes the model’s vocabulary size. This study employs the OV circuit (Elhage et al., 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand how source tokens generally influence output logits. The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $W_{E}W^{\mathbf{h}}_{V}W_{O}^{\mathbf{h}}W_{U}\in\mathbb{R}^{|V|\times|V|}$. Following the previous works (Elhage et al., 2021; Dar et al., 2023), this formulation omits layer normalisation.
@@ Appendix F GPT-2 Model Architecture
Before processing by self-attention and MLP processing, each token’s state is normalized to stabilise learning and reduce training time:
@@ Appendix G Path Patching Details
Figure 7: Path patching results for (Top) symbolic inputs and non-symbolic inputs (Bottom) in the middle-term corruption setup. Note that negative heads are positively influencing heads, as we replace corrupted activation with clean ones in our noising approach.
@@ Appendix H Localizing Mover Heads
Appendix I Non-symbolic Circuit Analysis
We present the results of our non-symbolic circuit analysis (Figure 9).
Figure 9: Comprehensive results of the non-symbolic circuit analysis. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. (e)-(h) Average attention patching weights across the batch: (e) weights for token position $[p]$ attending, and (f)-(h) weights for the last token position attending. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Appendix J Generalizability to Other Syllogisms
To assess the generalizability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 3), and the result is shown in Figure 10. Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises. The validity of these syllogisms is determined solely by their logical form, which is characterised by mood and figure combinations. The mood of a syllogism is defined by the arrangement of four proposition types ("All (A)", "No (E)", "Some (I)", "Some … not (O)") across its two premises and conclusion. Conversely, the figure of a categorical syllogism is determined by the structural arrangement of terms within the constituent sentences. This approach enables the evaluation of syntactic logic in isolation from contextual interference and belief-bias effects.
@@ Appendix I Non-symbolic Circuit Analysis
Model S BC BI
Group 1: Erratic Pattern
Pythia-70M 0.19 0.02 0.02
Pythia-160M 0.00 0.12 0.21
Group 2: Compatible Pattern
Pythia-410M 0.99 0.40 0.51
Pythia-1B 0.30 0.31 0.43
Llama-3.2-1B 1.00 0.81 0.70
Group 3: Variant Pattern
Qwen2.5-0.5B 1.00 0.96 0.90
Qwen2.5-1.5B 1.00 0.99 0.90
Table 5: Accuracy comparison of different models across three evaluation categories: Symbolic (S), Belief-Consistent (BC), and Belief-Inconsistent (BI). Group 1 models exhibit erratic activation patterns, failing to establish a stable reasoning mechanism. Group 2 models demonstrate a compatible reasoning circuit with consistent information flow and suppression. Group 3, particularly the Qwen models, exhibit a functional but distinct suppression mechanism, where suppression is applied at the final token rather than the $[p]$ position.
Appendix K Generalizability Across Model Sizes
@@
Figure 11: Comprehensive results of the symbolic circuit analysis across different model sizes. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Appendix L Generalizability Across Different Models
Figures 12 and 13 represent the results of applying the same circuit analysis in different model families. The $m$-suppression heads for each model were manually selected based on their likelihood, determined by patching scores and spatial positioning within the activation pattern. The corresponding model accuracies in the three datasets (Symbolic, Belief-Consistent, and Belief-Inconsistent) are presented in Table 5.
Figure 12: Comprehensive results of the symbolic circuit analysis across different series of models (Pythia). (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Figure 13: Comprehensive results of the symbolic circuit analysis across different series of models (Qwen2.5, Llama3.2 and fine-tuned GPT-2 medium). (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Appendix M Reproducibility Statement
Our study is conducted using TransformerLens (Nanda and Bloom, 2022) on an Nvidia A100 GPU with 80GB of memory. The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous_url.com.
Appendix N Statistical Significance
For all experiments, we use 90 samples each for both symbolic and non-symbolic arguments to balance analytical depth with computational efficiency. This sample size sufficiently yields statistically significant activation patching results ($p<0.05$).

### v3

@@ 4.3 Generalization to Model Sizes
Next, we expand our analysis to different sizes of GPT-2 (small, large and XL). Here, the intervention results show similar suppression mechanism patterns and information flow across all models. However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix J for additional details). Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex. We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (see Figure 6(a)–(b), Appendix E).
4.4 Generalization to Model Families
@@
We further extend our investigation to assess whether a similar reasoning mechanism can be identified across different LMs. Specifically, we evaluate the circuit on: Pythia with 70M, 160M, 410M, and 1B parameters (Biderman et al., 2023); Qwen 2.5 with 0.5B and 1.5B parameters (Yang et al., 2024); Llama 3.2 with 1B parameters (et al., 2024); and additionally, GPT-2-medium after being fine-tuned on argumentative texts for critical thinking (Betz et al., 2021).
@@
Our results reveal distinct grouping patterns. In relatively small models—such as Pythia 70M and 160M—we observe erratic activation patterns that mirror the limitations seen in under-scaled models like GPT-2 small. In contrast, models with more than 0.5B parameters (namely, Pythia 410M, Pythia 1B, and Llama 3.2 1B) exhibit a compatible pattern in terms of information flow and suppressive mechanisms. Qwen exhibits a variant suppressive mechanism, with the suppression head operating at the last token position rather than $[p]$. Meanwhile, the fine-tuned GPT-2 medium model yields activation patterns that are nearly indistinguishable from its pre-trained counterpart, suggesting that the core reasoning circuit originates from pre-training rather than fine-tuning. Table 3 summarizes the accuracy of each model across the Symbolic, Belief-Consistent, and Belief-Inconsistent datasets (see further circuit analysis details in Appendix K).
Overall, our results indicate that while the reasoning circuit generalizes effectively across sufficiently scaled models, subtle architectural nuances can lead to distinct but comparable dynamics.
Model S BC BI
Group 1: Erratic Pattern
Pythia-70M 0.19 0.02 0.02
Pythia-160M 0.00 0.12 0.21
Group 2: Compatible Pattern
Pythia-410M 0.99 0.40 0.51
Pythia-1B 0.30 0.31 0.43
Llama-3.2-1B 1.00 0.81 0.70
Group 3: Variant Pattern
Qwen2.5-0.5B 1.00 0.96 0.90
Qwen2.5-1.5B 1.00 0.99 0.90
Table 3: Accuracy comparison across different models: symbolic (S), belief-consistent (BC), and inconsistent (BI). Group 1 models exhibit erratic patterns, failing to establish a stable reasoning mechanism. Group 2 models exhibit a compatible reasoning circuit with consistent information flow and suppression. Group 3 exhibits a functional but distinct suppression mechanism, where suppression is applied at the last token position.
@@ 6 Conclusion
In this study, we presented a comprehensive mechanistic interpretation of syllogistic reasoning in language models. By combining activation patching, embedding space analysis, and circuit ablation, we uncovered a structured error-correction mechanism characterized by the suppression of duplicated middle terms and the propagation of relevant information via mover heads. Our findings show that this circuit is necessary and sufficient for symbolic syllogistic reasoning. Moreover, we demonstrated that, while this reasoning mechanism generalizes across syllogistic schemes, model architectures and sizes, it remains susceptible to contamination from belief biases when instantiated with natural language inputs. Overall, our findings suggest that LMs learn transferable reasoning mechanisms but that, at the same time, such mechanisms might be contaminated and suppressed by the same world knowledge acquired during pre-training. Further studies will be required to understand how the discovered symbolic circuit can be effectively disentangled from world knowledge to enable systematic generalization in LMs’ reasoning.
@@ 7 Limitations
It is important to acknowledge some of the limitations of our study. First, our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios which are hard to model via causal intervention techniques. Second, Our experimentation is confined to specific syllogistic reasoning templates due to the intricate level of granularity and the computational complexity required in circuit analysis. Despite these limitations, however, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by language models for formal reasoning and related biases, laying a solid foundation for future research in this field.
@@ Appendix A Unconditionally Valid Syllogism Schemes
Table 4: 15 Unconditionally valid syllogism schemes. The table lists the syllogisms by their traditional names, moods, and figures, with formalized logical expressions on the first line and corresponding natural language examples on the second line. The minor premise (m) is presented before the major premise (M) to emphasize the transitive property in AAA-1 syllogism.
@@ B.2 Non-symbolic
  • Constraints: We exclude terms tokenized into multiple tokens to maintain consistency in comparison with the symbolic dataset, which is essential for coherent circuit analysis.
  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI, 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent. This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
@@ B.3 Data Statistics
For all experiments, we use 90 samples each for both symbolic and non-symbolic arguments to balance analytical depth with computational efficiency. This sample size sufficiently yields statistically significant activation patching results ($p<0.05$). We organize data statistics for generated datasets (Table 5).
@@ Appendix C Embedding Space Analysis Details
One established method for analyzing hidden activation in transformer language models is the logit lens, which projects activation into embedding space (Nostalgebraist, 2020; Elhage et al., 2021; Geva et al., 2022). We focus on an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024), utilizing both unembedding ($W_{U}$) and embedding ($W_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ denotes the model’s vocabulary size. This study employs the OV circuit (Elhage et al., 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand how source tokens generally influence output logits. The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalized as $W_{E}W^{\mathbf{h}}_{V}W_{O}^{\mathbf{h}}W_{U}\in\mathbb{R}^{|V|\times|V|}$. Following the previous works (Elhage et al., 2021; Dar et al., 2023), this formulation omits layer normalization.
@@ Appendix F GPT-2 Model Architecture
Before processing by self-attention and MLP processing, each token’s state is normalized to stabilize learning and reduce training time:
@@ Appendix G Path Patching Details
Figure 7: Path patching results for (Top) symbolic inputs and non-symbolic inputs (Bottom) in the middle-term corruption setup. Note that negative heads are positively influencing heads, as we replace corrupted activation with clean ones in our noising method.
@@ Appendix H Localizing Mover Heads
Appendix I Generalizability to Other Syllogisms
To assess the generalizability of our model, we applied the circuit ablation method to all 15 unconditionally valid syllogisms (Table 4), and the result is shown in Figure 9. Unconditionally valid syllogisms are logical arguments that maintain their validity irrespective of the truth values of their premises. The validity of these syllogisms is determined solely by their logical form, which is characterized by mood and figure combinations. The mood of a syllogism is defined by the arrangement of four proposition types ("All (A)", "No (E)", "Some (I)", "Some … not (O)") across its two premises and conclusion. Conversely, the figure of a categorical syllogism is determined by the structural arrangement of terms within the constituent sentences. This approach enables the evaluation of syntactic logic in isolation from contextual interference and belief-bias effects.
@@ Appendix I Generalizability to Other Syllogisms
Appendix J Generalizability Across Model Sizes
@@
Figure 10: Comprehensive results of the symbolic circuit analysis across different model sizes. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, a dash (–) indicates the averaged values for tokens appearing between terms.
Appendix K Generalizability Across Different Models
Figures 11 and 12 represent the results of applying the same circuit analysis in different model families. The $m$-suppression heads for each model were manually selected based on their likelihood, determined by patching scores and positioning within the activation pattern.
Figure 11: Comprehensive results of the symbolic circuit analysis across different series of models (Pythia). (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, a dash (–) indicates the averaged values for tokens appearing between terms.
Figure 12: Comprehensive results of the symbolic circuit analysis across different series of models (Qwen2.5, Llama3.2 and fine-tuned GPT-2 medium). (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $m$-suppression head, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. For clarity, a dash (–) indicates the averaged values for tokens appearing between terms.